# PEARL: Preconditioner Enhancement through Actor-critic Reinforcement Learning

### v1

@@ Abstract
We present PEARL (Preconditioner Enhancement through Actor-critic Reinforcement Learning), a novel approach to learning matrix preconditioners. Existing preconditioners such as Jacobi, Incomplete LU, and Algebraic Multigrid methods offer problem-specific advantages but rely heavily on hyperparameter tuning. Recent advances have explored using deep neural networks to learn preconditioners, though challenges such as misbehaved objective functions and costly training procedures remain. PEARL introduces a reinforcement learning approach for learning preconditioners, specifically, a contextual bandit formulation. The framework utilizes an actor-critic model, where the actor generates the incomplete Cholesky decomposition of preconditioners, and the critic evaluates them based on reward-specific feedback. To further guide the training, we design a dual-objective function, combining updates from the critic and condition number. PEARL contributes a generalizable preconditioner learning method, dynamic sparsity exploration, and cosine schedulers for improved stability and exploratory power. We compare our approach to traditional and neural preconditioners, demonstrating improved flexibility and iterative solving speed.
@@ 1 Introduction
Many problems involving complex partial differential equations (PDEs) result in the discretization of space, producing large systems of linear equations. Solving these systems often requires advanced iterative solvers due to their size and complexity. Iterative solvers are foundational tools utilized in tasks such as numerical weather prediction, power system analysis, and fluid dynamics simulations. The behavior of these solvers is well-supported by a rich body of literature, much of which focuses on optimization techniques. This papers focuses is on the well-established domain of preconditioning [1].
@@ 2 Literature Review
Recent works utilize constrained optimization for three core techniques: learning to approximate the non-zero entries of traditional methods [7], optimizing for speedups of traditional methods [8], and predicting direct inverse preconditioners [9]. Other approaches utilizing fourier neural operators (FNOs) [10] and physics-informed neural networks (PINNs)[11], can further constrain the system based on the properties of the PDEs. Our approach aims to develop a general-purpose algorithm, similar to Chen’s work [12], for directly predicting inverse preconditioners. We extend this large body of work by employing reinforcement learning, allowing feedback from iterative solvers to indirectly backpropagate through the model during training [13].
@@ 3.2 Contextual Bandit
$pi^{*}=\arg\max_{\pi}\mathbb{E}\left[\sum_{t=1}^{T}r_{t}\right].$ (4)
@@ 4.1 Training Data Generation
These resulting systems are symmetric and positive definite. Their most prominent feature is extreme ill-conditioning. The 2D diffusion equation, particularly for large values of $h_{i}$, is notorious for producing highly ill-conditioned matrices. When combined with random perturbations, empirical results show that solving these systems, utilizing no preconditioner, becomes intractable. Although the random perturbations detract from the model’s physical interpretability, they introduce additional complexity and facilitate the mapping of a class of ill-conditioned systems to a set of effective preconditioning techniques. Our experiments utilize a $25\times 25$ discretization, resulting in condition numbers typically ranging from 3,000 to 12,000.
@@ 4.5 Training Procedure
To train both the actor and the critic, we utilize an experience replay mechanism. We first generate perturbed systems and then apply two forms of exploration noise: symmetric and sparse. Symmetric noise is defined as standard noise sampled from $N(0,\Sigma)$ with a size of $\mathbb{R}^{n\times n}$. The noise is applied as follows:
@@ 7 Discussion
Using this framework to learn the ILU decomposition instead of the incomplete Cholesky decomposition, we observe that the model significantly reduces the condition number but fails to accelerate the solver. This outcome arises because the LU decomposition does not inherently preserve the symmetric positive definite (SPD) property. Consequently, the condition gradient becomes unstable, resulting in suboptimal solutions, as illustrated in Figure 6. By incorporating the critic into the optimization process, we mitigate collapse into suboptimal solutions; however, we still encounter challenges in fitting an effective preconditioner. This observation leads us to hypothesize that the condition number primarily guides the actor toward the optimal solution, while the critic facilitates structured exploration of the search space through gradual, incremental updates. This finding is significant as it suggests that this reinforcement learning framework can be further analyzed to quantify the balance between exploitation and exploration in the learned preconditioner policies. Additionally, we provide outputs from timesteps 2500, 1000, and 100 of the critic-only cosine-scheduled model, shown at the bottom of Figure 4.
Limitations. While the proposed framework is evaluated on 2D diffusion equations with random perturbations—thereby capturing a class of ill-conditioned systems—its performance on more diverse PDEs, particularly those involving nonlinearities, complex boundary conditions, or domain-specific structures, remains untested. Furthermore, although the experimental setup includes moderate-size matrices, large-scale scientific applications involve systems with millions of degrees of freedom. Training an actor-critic agent on matrices of this size may require extensive memory, complex sampling strategies, and distributed computing infrastructures. Assessing whether the method can be feasibly scaled up without sacrificing accuracy or incurring prohibitive computational costs is an open question. Plus, the training procedure uses both condition-number gradients and rewards derived from convergence metrics, each of which can be numerically delicate. Condition-number calculations can become unstable for matrices approaching singularity, leading to noisy gradient signals. Although the introduction of a cosine scheduler helps moderate these effects, the stability of the training loop still hinges on heuristics such as scheduler periods, gradient clipping thresholds, and loss-function weights. An inadequate choice of hyperparameters risks suboptimal convergence or unstable training dynamics.

### v2

@@ Abstract
We present PEARL (Preconditioner Enhancement through Actor-critic Reinforcement Learning), a novel approach to learning matrix preconditioners. Existing preconditioners such as Jacobi, Incomplete LU, and Algebraic Multigrid methods offer problem-specific advantages but rely heavily on hyperparameter tuning. Recent advances have explored using deep neural networks to learn preconditioners, though challenges such as misbehaved objective functions and costly training procedures remain. PEARL introduces a reinforcement learning approach for learning preconditioners, specifically, a contextual bandit formulation. The framework uses an actor-critic model, where the actor generates the incomplete Cholesky decomposition of preconditioners, and the critic evaluates them based on reward-specific feedback. To further guide the training, we design a dual-objective function, combining updates from the critic and condition number. PEARL contributes a generalizable preconditioner learning method, dynamic sparsity exploration, and cosine schedulers for improved stability and exploratory power. We compare our approach to traditional and neural preconditioners, demonstrating improved flexibility and iterative solving speed.
@@ 1 Introduction
Many problems involving complex partial differential equations (PDEs) result in the discretization of space, producing large systems of linear equations. Solving these systems often requires advanced iterative solvers due to their size and complexity. Iterative solvers are foundational tools used in tasks such as numerical weather prediction, power system analysis, and fluid dynamics simulations. The behavior of these solvers is well-supported by a rich body of literature, much of which focuses on optimization techniques. This papers focuses is on the well-established domain of preconditioning [1].
@@ 2 Literature Review
Recent works use constrained optimization for three core techniques: learning to approximate the non-zero entries of traditional methods [7], optimizing for speedups of traditional methods [8], and predicting direct inverse preconditioners [9]. Other approaches use fourier neural operators (FNOs) [10] and physics-informed neural networks (PINNs)[11], can further constrain the system based on the properties of the PDEs. Our approach aims to develop a general-purpose algorithm, similar to Chen’s work [12], for directly predicting inverse preconditioners. We extend this large body of work by employing reinforcement learning, allowing feedback from iterative solvers to indirectly backpropagate through the model during training [13].
@@ 3.2 Contextual Bandit
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\left[\sum_{t=1}^{T}r_{t}\right].$ (4)
@@ 4.1 Training Data Generation
These resulting systems are symmetric and positive definite. Their most prominent feature is extreme ill-conditioning. The 2D diffusion equation, particularly for large values of $h_{i}$, is notorious for producing highly ill-conditioned matrices. When combined with random perturbations, empirical results show that solving these systems, utilizing no preconditioner, becomes intractable. Although the random perturbations detract from the model’s physical interpretability, they introduce additional complexity and facilitate the mapping of a class of ill-conditioned systems to a set of effective preconditioning techniques. Our experiments use a $25\times 25$ discretization, resulting in condition numbers typically ranging from 3,000 to 12,000.
@@ 4.5 Training Procedure
To train both the actor and the critic, we use an experience replay mechanism. We first generate perturbed systems and then apply two forms of exploration noise: symmetric and sparse. Symmetric noise is defined as standard noise sampled from $N(0,\Sigma)$ with a size of $\mathbb{R}^{n\times n}$. The noise is applied as follows:
@@ 7 Discussion
Using this framework to learn the ILU decomposition instead of the incomplete Cholesky decomposition, we observe that the model significantly reduces the condition number but fails to accelerate the solver. This outcome arises because the LU decomposition does not inherently preserve the symmetric positive definite (SPD) property. Consequently, the condition gradient becomes unstable, resulting in suboptimal solutions, as illustrated in Figure 6. By incorporating the critic into the optimization process, we mitigate collapse into suboptimal solutions; however, we still encounter challenges in fitting an effective preconditioner. This observation leads us to hypothesize that the condition number primarily guides the actor toward the optimal solution, while the critic facilitates structured exploration of the search space through gradual, incremental updates. This finding is significant as it suggests that this reinforcement learning framework can be further analyzed to quantify the balance between exploitation and exploration in the learned preconditioner policies. Furthermore, we provide outputs from time steps 2500, 1000, and 100 of the critic-only cosine-scheduled model, shown at the bottom of Figure 4.
Limitations. While the proposed framework is evaluated on 2D diffusion equations with random perturbations—thereby capturing a class of ill-conditioned systems—its performance on more diverse PDEs, particularly those involving nonlinearities, complex boundary conditions, or domain-specific structures, remains untested. Furthermore, although the experimental setup includes moderate-size matrices, large-scale scientific applications involve systems with millions of degrees of freedom. Training an actor-critic agent on matrices of this size may require extensive memory, complex sampling strategies, and distributed computing infrastructures. Assessing whether the method can be feasibly scaled up without sacrificing accuracy or incurring prohibitive computational costs is an open question. In addition, the training procedure uses both condition-number gradients and rewards derived from convergence metrics, each of which can be numerically delicate. Condition-number calculations can become unstable for matrices approaching singularity, leading to noisy gradient signals. Although the introduction of a cosine scheduler helps moderate these effects, the stability of the training loop still depends on heuristics such as scheduler periods, gradient clipping thresholds, and loss-function weights. An inadequate choice of hyperparameters risks suboptimal convergence or unstable training dynamics.